# Stable coresets: Unleashing the power of uniform sampling

### v1

@@ Abstract
Uniform sampling is a highly efficient method for data summarization. However, its effectiveness in producing coresets for clustering problems is not yet well understood, primarily because it generally does not yield a strong coreset, which is the prevailing notion in the literature. We formulate stable coresets, a notion that is intermediate between the standard notions of weak and strong coresets, and effectively combines the broad applicability of strong coresets with highly efficient constructions, through uniform sampling, of weak coresets. Our main result is that a uniform sample of size $O(\epsilon^{-2}\log d)$ yields, with high constant probability, a stable coreset for $1$-median in $\mathbb{R}^{d}$ under the $\ell_{1}$ metric. We then leverage the powerful properties of stable coresets to easily derive new coreset constructions, all through uniform sampling, for $\ell_{1}$ and related metrics, such as Kendall-tau and Jaccard. We also show applications to fair clustering and to approximation algorithms for $k$-median problems in these metric spaces. Our experiments validate the benefits of stable coresets in practice, in terms of both construction time and approximation quality.
@@ 1 Introduction
However, the advantages of strong coresets come with nontrivial computational challenges, as their construction algorithm must inevitably read the entire dataset.  [^1] Consider $k=1$ and a dataset where all points are densely clustered, except for one ”outlier” point extremely far away. While this outlier hardly affects the optimal center, its impact on the objective value might be significant, and thus it must be examined and even included in every strong coreset. Can we construct coresets in sublinear time? Weak coresets demonstrate that this is possible, as they can sometimes be constructed through uniform sampling [HJL23, HJLW23, MF19], a highly efficient method that is easy to implement in streaming and distributed settings, and is well-known to be extremely useful in practice. However, uniform sampling cannot reliably capture small clusters, which motivates us to study the fundamental case $k=1$, where uniform sampling can succeed without additional assumptions about the dataset. This case arises naturally in many practical scenarios and serves as an important building block for the general problem.
@@ 1.1 Problem setup and definitions
While not formalized as a coreset notion, the principle underlying (3) and its compatibility with uniform sampling were first used by Indyk [Ind99, Ind01], to compare the costs of two centers in the context of $1$-median with discrete centers (i.e., the center must be one of the dataset points). For finite $\mathcal{X}$, Indyk’s analysis would yield a stable $(0,\epsilon)$-coreset of size $O(\epsilon^{-2}\log|\mathcal{X}|)$ (details in Appendix A.1).
@@ 1.2 Our contribution
Let $P\subset\mathbb{R}^{d}$ be finite and let $0<\epsilon\leq\frac{1}{5}$. Then, a uniform sample of size $O(\epsilon^{-2}\log d)$ from $P$ is a stable $(\epsilon/6,4\epsilon)$-coreset for $1$-median in $\ell_{1}^{d}$ with probability at least $4/5$.
@@ 1.4 Related work
In $\ell_{p}$ metric spaces, the $k$-median problem for general $k$ is APX-hard [GI03, Tre00], with some recent advances about its inapproximability [CKL22]. In a metric space of bounded doubling dimension $D$, this problem admits a polynomial-time approximation scheme (PTAS), namely, $(1+\epsilon)$-approximation that runs in time $\tilde{O}(2^{(\frac{1}{\epsilon})^{O(D^{2})}}n)$ [CFS21]. Clearly this approach is only practical when the doubling dimensions is very low.
@@
In comparison, strong coresets for $k$-median in $\ell_{1}$ of size $\operatorname{poly}(k/\epsilon)$ follow implicitly from [JKLZ24], because $\ell_{1}$ is contained in $\ell_{2}$-squared, however, constructing such coresets requires at least linear time.
Coresets for $1$-center (aka Minimum Enclosing Ball) in $\ell_{1}$ and in related metrics were studied in [CGJK25]. It was shown that for both strong and weak coresets, the coreset size must depend on the dimension (in contrast to $1$-median).
@@ 2 Preliminaries
    Every strong $\epsilon$-coreset of $P$, for $0<\epsilon\leq\frac{1}{5}$, is also a stable $(\epsilon,4\epsilon)$-coreset.
@@
    for every $P\subseteq\mathcal{X}_{1}$ and $c\in P$, $\operatorname{cost}(c,P)=\operatorname{cost}(f(c),f(P))$.
@@ 3 A framework for stable coresets
Let $P\subseteq\mathcal{X}$ and $0<\epsilon\leq\frac{1}{5}$, and suppose $\operatorname{\overline{cost}}(\mu,Q)\leq c\cdot\operatorname{\overline{cost}}(\mu,P)$ for some $c\geq 1$. If $Q\subseteq P$ is an $\epsilon$-RCDA of $P$ in $\mathcal{X}$ then $Q$ is a $(\frac{\epsilon}{c},4\epsilon)$-stable coreset of $P$.
@@ 4 Stable coresets in $\ell_{1}$ through uniform sampling
We now bound the VC dimension of the class of threshold functions $\mathcal{T}$, showing that it essentially logarithmic in the dimension $d$ (see also [Gey18] which provides tight bounds).
@@
Using a theorem established by Yi, Long and Srinivasan [LLS01], we can bound the size of such $\epsilon$-approximation (see also [HS11]).
@@
Let $\mathcal{F}$ be a class of function from $P$ to $\{0,1\}$, with finite VC dimension, and let $\mathcal{D}$ be some probability distribution over $P$. Then, with probability at least $1-\delta$, a random sample $Q\subseteq P$ of size $O\left(\epsilon^{-2}(\operatorname{VCdim}(\mathcal{F})+\log\frac{1}{\delta})\right)$ satisfies
@@
Let $\epsilon\in(0,1)$. If $Q$ is an $\epsilon$-approximation of $P$, then $Q$ is a $20\epsilon$-RCDA of $P$.
@@ 5 Applications
Let $f:\mathcal{X}_{1}\rightarrow\mathcal{X}_{2}$ be an embedding between metric spaces $(\mathcal{X}_{1},\text{dist}_{1})$ and $(\mathcal{X}_{2},\text{dist}_{2})$ with distortion $D$. For every $Q\subseteq P\subseteq\mathcal{X}_{1}$, if $f(Q)$ is a stable $(\epsilon,\eta)$-coreset of $f(P)$ in $\mathcal{X}_{2}$ for some $\epsilon,\eta>0$, and the values $\epsilon^{\prime}:=(1+\epsilon)/D^{2}-1$ and $\eta^{\prime}:=D^{2}(1+\eta)-1$ are positive, then $Q$ is a stable $(\epsilon^{\prime},\eta^{\prime})$-coreset of $P$ in $\mathcal{X}_{1}$.
This proposition extends our results in Theorem 1.4 to metric spaces that can be embedded into $\ell_{1}$ with small distortion. For example, our results extend to the Euclidean metric using Dvoretzky’s Theorem, yielding stable coresets of size $O(\epsilon^{-2}\log(d/\epsilon))$ for $1$-median in $\ell_{2}^{d}$, see Section D.2. We remark that it suffices to have the distortion guarantee (5) only for pairs that involve a point from $P$, which is known in the literature as terminal embedding, see Section D.1.
@@
In metric spaces that admit stable coresets through uniform sampling, we can employ the framework introduced by Kumar, Sabharwal, and Sen [KSS04] and further refined for additional metric spaces by Ackermann, Blömer and Sohler [ABS10]. These papers design approximation algorithms for $k$-median in metric spaces that have the property that for every instance $P$, a $1$-median of $P$ is approximated (with high probability) by an optimal, or approximately optimal, solution for a uniform sample $Q\subseteq P$. We restate the main theorem in [ABS10] using stable coresets, and it can now be applied to several metrics where it was previously unknown, including Hamming, Kendall-tau and Jaccard, thereby extending existing $1$-median algorithms to the more general $k$-median problem.
Theorem 5.4 ([ABS10], Theorem 1.1).
@@
Let $P\subset\mathbb{R}^{d}$ be finite and $C$-dispersed, and let $\epsilon\in(0,\frac{1}{5})$. Then a uniform sample of size $O(C^{2}\epsilon^{-2}\log d)$ from $P$ is a strong $\epsilon$-coreset for $1$-median in $\ell_{1}^{d}$ with probability at least $\frac{3}{4}$.
@@ 6 Experiments
All experiments were conducted on a PC with Apple M1 and 16GB RAM running Python 3.9.6 on Darwin 22.6.0. For each experiment, we report the average results over 20 independent runs to ensure statistical significance. The datasets used in our experiments are detailed in Table 1. The source code used to run the experiments is available at anonymous.4open.science/r/StableCoresets-A8CE.
@@
Yellow Taxi NYC (YT) [NYC24] 2.8MM 11 New York City taxi trips in Jan. 2024
Twitter [CGS18] 1.3MM 3 Timestamp, latitude, longitude of tweets
Single-Cell Gene Expression (SCGE) [10x19] 7,865 33,586 Peripheral blood mononuclear cells gene expression
@@
We compare uniform sampling against the importance sampling-based coreset construction proposed in [JKLZ24]. Their method works by iteratively computing sensitivity scores for each point and sampling progressively smaller subsets, with each iteration reducing the size by a logarithmic factor until achieving a dimension-independent coreset. In our implementation, we evaluated their approach using both one and two iterations of this reduction process. The comparison focuses on the $1$-median problem in $\ell_{1}$ metric across three datasets from Table 1; Yellow Taxi NYC, Twitter and Single-Cell Gene Expression.
@@
Using the MAL dataset, we aggregated rankings of $234K$ users over $50$ anime titles. We implemented five widely used rank aggregation methods: three Markov Chain-based approaches (MC1, MC2, MC3), Borda’s sorting algorithm and scaled footrule aggregation (SFO) [DKNS01, KSSJ17]. Figure 2 illustrates the relative error of solutions computed on coresets of different sizes. It is important to note that these heuristics do not directly optimize the Kendall-tau cost. Consequently, solutions computed on the coresets can occasionally yield lower Kendall-tau costs than those from the original dataset, resulting in negative relative error values. The results confirm that relatively small coresets achieve results comparable to those obtained on the original dataset, even when using heuristic approaches.
Figure 2: Comparison of ranking method performance with respect to coreset sizes. The plot shows relative error (%) between coreset approximation and original dataset results. Regression lines demonstrate error trends as coreset size increases, with data points marking actual experimental measurements.
@@
In this experiment, we sampled coresets of varying sizes from the dataset and then applied fairness constraints based on an arbitrary partitioning of items into two groups. We implemented the fairness-constrained integer linear programming algorithm by Kuhlman and Rundensteiner [KR20] on both the original dataset and the coresets. Due to algorithmic constraints, we restricted the dataset to 8,700 user rankings, with each user ranking 16 anime titles. This algorithm optimizes the Kendall-tau cost objective while enforcing the fairness measure as a linear constraint. Negative error values occur because the ILP solver might find slightly better solutions on the coreset rather than on the original dataset. Figure 3 demonstrates that solutions obtained from even-small sized coresets closely approximate those from the original dataset. This confirms that uniform sampling produces stable coresets that effectively support constraint-based optimization, even when those constraints were not considered during the sampling process.
Figure 3: Impact of fairness constraints on coreset approximation error. The parameter $p$ represents the probability of sampling popular anime titles across two distinct groups. $p=0.5$ indicates balanced sampling between groups, while lower $p$ values indicate one group containing predominantly less popular anime. Shaded regions represent the standard error of the mean across multiple experimental runs.
@@
For this we utilized the high-dimensional Single-Cell Gene Expression dataset (7,865 samples across 33,586 dimensions). For a given dimension count $d$, we randomly selected $d$ dimensions from the input dataset and then uniformly sampled a coreset (for various coreset size 150,300,500). For each coreset we measured the relative error as defined in Equation 6. Figure 4 illustrates the relationship between dimension count and error rates. We note that the slight upward trend in error is very slight and likely attributable to sampling variance rather than dimensional dependency. This demonstrate that the relative error remains stable as dimension count increases and supports our conjecture that the theoretical bounds could be tightened further, and dimension-independent bound can be derived.
Figure 4: Relative error (%) versus number of dimensions (log scale) for different coreset sizes. Regression lines demonstrate error trends as the number of dimensions increases. The slope is outlined for each regression line. Note that as the coreset size increases, the variance decreases, making the slope coefficient more statistically meaningful.
@@ A.1 Stable coresets in finite metric spaces
We remind the reader the following result by Indyk [Ind01].
Theorem A.1 ([Ind01], Theorem 31).
@@ Appendix B Proofs omitted from Section 3
Plugging this into (7) and using our assumption that $\epsilon\leq\frac{1}{5}$, we conclude that

### v2

@@ Abstract
Uniform sampling is a highly efficient method for data summarization. However, its effectiveness in producing coresets for clustering problems is not yet well understood, primarily because it generally does not yield a strong coreset, which is the prevailing notion in the literature. We formulate stable coresets, a notion that is intermediate between the standard notions of weak and strong coresets, and effectively combines the broad applicability of strong coresets with highly efficient constructions, through uniform sampling, of weak coresets. Our main result is that a uniform sample of size $O(\epsilon^{-2}\log d)$ yields, with high constant probability, a stable coreset for $1$-median in $\mathbb{R}^{d}$ under the $\ell_{1}$ metric. We then leverage the powerful properties of stable coresets to easily derive new coreset constructions, all through uniform sampling, for $\ell_{1}$ and related metrics, such as Kendall-tau and Jaccard. We also show applications to fair rank aggregation and to approximation algorithms for $k$-median problem in these metric spaces. Our experiments validate the benefits of stable coresets in practice, in terms of both construction time and approximation quality.
@@ 1 Introduction
However, the advantages of strong coresets come with nontrivial computational challenges, as their construction algorithms must inevitably read the entire dataset. [^1] Consider $k=1$ and a dataset where all points are densely clustered, except for one ”outlier” point extremely far away. While this outlier hardly affects the optimal center, its impact on the objective value might be significant, and thus it must be examined and even included in every strong coreset. Can we construct coresets in sublinear time? Weak coresets demonstrate that this is possible, as they can sometimes be constructed through uniform sampling [HJL23, HJLW23, MF19], a highly efficient method that is easy to implement in streaming and distributed settings, and is well-known to be extremely useful in practice. However, uniform sampling cannot reliably capture small clusters, which motivates us to study the fundamental case $k=1$, where uniform sampling can succeed without additional assumptions about the dataset. This case arises naturally in many practical scenarios and serves as an important building block for the general problem.
@@ 1.1 Problem setup and definitions
While not formalized as a coreset notion, the principle underlying (3) and its compatibility with uniform sampling were first used by [Ind99, Ind01], to compare the costs of two centers in the context of $1$-median with discrete centers (i.e., the center must be one of the dataset points). For finite $\mathcal{X}$, Indyk’s analysis would yield a stable $(0,\epsilon)$-coreset of size $O(\epsilon^{-2}\log|\mathcal{X}|)$ (details in Appendix A.1).
@@ 1.2 Our contribution
Let $P\subset\mathbb{R}^{d}$ be finite and let $\epsilon\in(0,\frac{1}{5})$. Then, a uniform sample of size $O(\epsilon^{-2}\log d)$ from $P$ is a stable $(\epsilon/6,4\epsilon)$-coreset for $1$-median in $\ell_{1}^{d}$ with probability at least $4/5$.
@@ 1.4 Related work
In $\ell_{p}$ metric spaces, the $k$-median problem for general $k$ is APX-hard [GI03, Tre00], with some recent advances about its inapproximability [CKL22]. In a metric space of bounded doubling dimension $D$, this problem admits a polynomial-time approximation scheme (PTAS), namely, $(1+\epsilon)$-approximation that runs in time $\tilde{O}(2^{(\frac{1}{\epsilon})^{O(D^{2})}}n)$ [CFS21]. Clearly this approach is only practical when the doubling dimension is very small.
@@
In comparison, strong coresets for $k$-median in $\ell_{1}$ of size $\operatorname{poly}(k/\epsilon)$ follow implicitly by [JKLZ24], because $\ell_{1}$ is contained in $\ell_{2}$-squared, however, constructing such coresets requires at least linear time.
Coresets for $1$-center (aka Minimum Enclosing Ball) in $\ell_{1}$ and in related metrics were studied by [CGJK25]. It was shown that for both strong and weak coresets, the coreset size must depend on the dimension (in contrast to $1$-median).
@@ 2 Preliminaries
    Every strong $\epsilon$-coreset of $P$, for $\epsilon\in(0,\frac{1}{5})$, is also a stable $(\epsilon,4\epsilon)$-coreset.
@@
    for every $P\subseteq\mathcal{X}_{1}$ and $c\in\mathcal{X}_{1}$, $\operatorname{cost}(c,P)=\operatorname{cost}(f(c),f(P))$.
@@ 3 A framework for stable coresets
Let $P\subseteq\mathcal{X}$ and $\epsilon\in(0,\frac{1}{5})$, and suppose $\operatorname{\overline{cost}}(\mu,Q)\leq c\cdot\operatorname{\overline{cost}}(\mu,P)$ for some $c\geq 1$. If $Q\subseteq P$ is an $\epsilon$-RCDA of $P$ in $\mathcal{X}$ then $Q$ is a $(\frac{\epsilon}{c},4\epsilon)$-stable coreset of $P$.
@@ 4 Stable coresets in $\ell_{1}$ through uniform sampling
We now bound the VC dimension of the class of threshold functions $\mathcal{T}$, showing that it is essentially logarithmic in the dimension $d$. While tight bounds appear in [Gey18], we provide a shorter proof for completeness.
@@
Using a theorem established by [LLS01], we can bound the size of such $\epsilon$-approximation (see also [HS11]).
@@
Let $\mathcal{F}$ be a class of function from $P$ to $\{0,1\}$, with finite VC dimension, and let $\mathcal{D}$ be some probability distribution over $P$. Then, with probability at least $1-\delta$, a random sample $Q\sim D$ of size $O\left(\epsilon^{-2}(\operatorname{VCdim}(\mathcal{F})+\log\frac{1}{\delta})\right)$ satisfies
@@
Let $\epsilon\in(0,1)$ and let $P$ be a finite set in $\ell_{1}^{d}$. If $Q$ is an $\epsilon$-approximation of $P$, then $Q$ is a $20\epsilon$-RCDA of $P$.
@@ 5 Applications
Let $f:\mathcal{X}_{1}\rightarrow\mathcal{X}_{2}$ be an embedding between metric spaces $(\mathcal{X}_{1},\text{dist}_{1})$ and $(\mathcal{X}_{2},\text{dist}_{2})$ with distortion $D$ and scaling factor $r$. For every $Q\subseteq P\subseteq\mathcal{X}_{1}$, if $f(Q)$ is a stable $(\epsilon,\eta)$-coreset of $f(P)$ in $\mathcal{X}_{2}$ for some $\epsilon,\eta>0$, and the values $\epsilon^{\prime}:=(1+\epsilon)/D^{2}-1$ and $\eta^{\prime}:=D^{2}(1+\eta)-1$ are positive, then $Q$ is a stable $(\epsilon^{\prime},\eta^{\prime})$-coreset of $P$ in $\mathcal{X}_{1}$.
This proposition extends our results in Theorem 1.4 to metric spaces that can be embedded into $\ell_{1}$ with small distortion. For example, our results extend to the Euclidean metric using Dvoretzky’s Theorem [Gor88, Sch06], yielding stable coresets of size $O(\epsilon^{-2}\log(d/\epsilon))$ for $1$-median in $\ell_{2}^{d}$, see Section D.2. We remark that it suffices to have the distortion guarantee (5) only for pairs that involve a point from $P$, which is known in the literature as terminal embedding, see Section D.1.
@@
In metric spaces that admit stable coresets through uniform sampling, we can employ the framework introduced by  [KSS04] and further refined for additional metric spaces by  [ABS10]. These papers design approximation algorithms for $k$-median in metric spaces that have the property that for every instance $P$, a $1$-median of $P$ is approximated (with high probability) by an optimal, or approximately optimal, solution for a uniform sample $Q\subseteq P$. We restate the main theorem in [ABS10] using the language of stable coresets, and it can now be applied to several metrics where it was previously unknown, including Hamming, Kendall-tau and Jaccard, thereby extending existing $1$-median algorithms to the more general $k$-median problem.
Theorem 5.4 (Theorem 1.1 in [ABS10]).
@@
Let $P\subset\mathbb{R}^{d}$ be finite and $C$-dispersed, and let $\epsilon\in(0,\frac{1}{5})$. Then a uniform sample of size $O(C\epsilon^{-2}\log\frac{d}{\delta})$ from $P$ is a strong $\epsilon$-coreset for $1$-median in $\ell_{1}^{d}$ with probability at least $1-\delta$.
@@ 6 Experiments
All experiments were conducted on a PC with Apple M1 and 16GB RAM running Python 3.9.6 on Darwin 22.6.0. For each experiment, we report the average results over 20 independent runs to ensure statistical significance. The datasets used in our experiments are detailed in Table 1. The source code used to run the experiments is available at https://github.com/amircarmel-lab/StableCoresets.
@@
Yellow Taxi NYC (YT) [NYC24] 2.8M 11 New York City taxi trips in Jan. 2024
Twitter [CGS18] 1.3M 3 Timestamp, latitude, longitude of tweets
Single-Cell Gene
Expression (SCGE) [10x19]
7,865 33,586 Peripheral blood mononuclear cells gene expression
@@
We compare uniform sampling against the importance sampling-based coreset construction proposed by [JKLZ24]. Their method works by iteratively computing sensitivity scores for each point and sampling progressively smaller subsets, with each iteration reducing the size by a logarithmic factor until achieving a dimension-independent coreset. In our implementation, we evaluated their approach using both one and two iterations of this reduction process. The comparison focuses on the $1$-median problem in $\ell_{1}$ metric across three datasets from Table 1; Yellow Taxi NYC, Twitter and Single-Cell Gene Expression.
@@
Using the MAL dataset, we aggregated rankings of $234K$ users over $50$ anime titles. We implemented five widely used rank aggregation methods: three Markov Chain-based approaches (MC1, MC2, MC3), Borda’s sorting algorithm and scaled footrule aggregation (SFO) [DKNS01, KSSJ17]. Figure 2(a) illustrates the relative error of solutions computed on coresets of different sizes. It is important to note that these heuristics do not directly optimize the Kendall-tau cost. Consequently, solutions computed on the coresets can occasionally yield lower Kendall-tau costs than those from the original dataset, resulting in negative relative error values. The results confirm that relatively small coresets achieve results comparable to those obtained on the original dataset, even when using heuristic approaches.
(a) Comparison of ranking method performance with respect to coreset sizes. The plot shows relative error (%) between coreset approximation and original dataset results. Regression lines demonstrate error trends as coreset size increases, with data points marking actual experimental measurements.
(b) Impact of fairness constraints on coreset approximation error. The parameter $p$ represents the probability of sampling popular anime titles across two distinct groups. $p=0.5$ indicates balanced sampling between groups, while lower $p$ values indicate one group containing predominantly less popular anime. Shaded regions represent the standard error of the mean across multiple experimental runs.
(c) Relative error (%) versus number of dimensions (log scale) for different coreset sizes. Regression lines demonstrate error trends as the number of dimensions increases. The slope is outlined for each regression line. Note that as the coreset size increases, the variance decreases, making the slope coefficient more statistically meaningful.
@@
In this experiment, we sampled coresets of varying sizes from the dataset and then applied fairness constraints based on an arbitrary partitioning of items into two groups. We implemented the fairness-constrained integer linear programming algorithm by [KR20] on both the original dataset and the coresets. Due to algorithmic constraints, we restricted the dataset to 8,700 user rankings, with each user ranking 16 anime titles. This algorithm optimizes the Kendall-tau cost objective while enforcing the fairness measure as a linear constraint. Negative error values occur because the ILP solver might find slightly better solutions on the coreset rather than on the original dataset. Figure 2(b) demonstrates that solutions obtained from even small-sized coresets closely approximate those from the original dataset. This confirms that uniform sampling produces stable coresets that effectively support constraint-based optimization, even when those constraints were not considered during the sampling process.
@@
To test this, we utilized the high-dimensional Single-Cell Gene Expression dataset (7,865 samples across 33,586 dimensions). For a given dimension count $d$, we randomly selected $d$ dimensions from the input dataset and then uniformly sampled a coreset for various sizes (150, 300, 500). For each coreset, we measured the relative error as defined in Equation 6.
Figure 2(c) illustrates the relationship between dimension count and error rates. We note that the slight upward trend in error is likely attributable to sampling variance rather than dimensional dependency. This demonstrates that the relative error remains stable as dimension count increases, supporting our conjecture that the theoretical bounds could be tightened to yield a dimension-independent bound.
@@ A.1 Stable coresets in finite metric spaces
We remind the reader the following result by [Ind01].
Theorem A.1 (Theorem 31 in [Ind01]).
@@ Appendix B Proofs omitted from Section 3
Plugging this into (7) and using our assumption that $\epsilon<\frac{1}{5}$, we conclude that